OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

hero (hero):

Using y = mx + b, input each point into the formula separately to produce the following equations: 1 = 2m + b 2 = -3m + b Isolate b in each equation to get: b = 1- 2m b = 2 + 3m Set b = b to get: 1 - 2m = 2 + 3m Add 2m to both sides; Subtract 2 from both sides to get: 1 - 2 = 3m + 2m - 1 = 5m Divide both sides by 5 to get -1/5 = m Input -1/5 for m into one of the original equations and solve for b b = 1 - 2(-1/5) b = 1 + 2/5 b = 5/5 + 2/5 b = 7/5 Therefore the equation of the line is y = -x/5 + 7/5 or y = (7 - x)/5

hero (hero):

By the way, you didn't post the equations below.